OpenStudy (anonymous):

factor the trinomial completely x^2+x-42

OpenStudy (anonymous):

we want two numbers that add to +1 and multiply to -42. +7 and -6 meet this criteria. (x+7)(x-6)=x^2+x-42

OpenStudy (anonymous):

Two numbers when that add to the 1 when multiplied give you: -42. So you know -6 + 7 = 1. So you have x^2-6x+7x-42 now factor into groups x(x-6) + 7(x-6) so you're answer is (x-6)(x+7) = x^2+x-42

OpenStudy (anonymous):

I realise this is the maths forum, but "so YOUR answer is"